A bicycle tire registers 90 psi (14.7 psi = 1 atm). What is that in torr?
Answer #1

14.7 psi = 1 atm

90 psi = 90 / 14.7 = 6.12 atm

we know,

1 atm = 760 torr

and

6.12 atm = 6.12 * 760 = 4653 torr = 4.65 * 10^3 torr (answer)